OpenStudy (anonymous):

Jaheed is making punch for a party. The recipe calls for 3.5 liters of fruit juice. Write a formula to help Jaheed determine the number of cartons of juice he needs to buy to make the punch. Let's let n = number of cartons of juice m = number of liters in each carton

OpenStudy (paki):

m * n = 3.5

OpenStudy (yanasidlinskiy):

m * n = 3.5 n cartons * m litres/carton = mn litres = 3.5L so you need n = 3.5/m cartons

OpenStudy (conqueror):

If each carton contained 0.5 liters you would need n = 3.5/0.5 = 7 cartons. So you should be able to write the formula now giving n in terms of m If the recipe needed 3.8 liters of juice then you would get n=3.6/0.5=7.2 cartons which you would have to round to 8 since 7 would not be enough. The formula is n=3.5/m rounded up to the nearest whole number

OpenStudy (anonymous):

ok im kinda understanding it a littel at least? next question i post ima try and solve it out and can you guys guide me???

OpenStudy (yanasidlinskiy):

Yes!

OpenStudy (paki):

yeah

OpenStudy (conqueror):

You need to know how much a carton holds.